# Assessing the replicability of RCTs in RWE emulations

### v1

@@ 3.1.1 The original formulation
The sceptical $p$-value can be interpreted quantitatively: Smaller $p_{S}$ values indicate a higher degree of replication success. The sceptical $p$-value is one-sided to ensure that replication can only occur if the replication effect estimate has the same sign as the original effect estimate. Further properties of the sceptical $p$-value are studied in Held, (2020). First, the approach can be shown to be more stringent than the two-trials rule: if replication success is achieved ($p_{S}\leq\alpha$), then two-trials rule is also fulfilled ($p_{\mbox{\scriptsize TTR}}=\max\{p_{\mbox{\scriptsize RCT}},p_{\mbox{\scriptsize RWE}}\}\leq\alpha$) but not necessarily the other way around. As a consequence, the overall Type-I error rate of the sceptical $p$-value (assuming that both the original and the replication effect estimate is zero) is smaller than the Type-I error rate $\alpha^{2}$ of the two-trials rule. Second, for fixed original and replication $p$-value, the sceptical $p$-value increases with increasing variance ratio $c$. The variance ratio can be rewritten as the relative sample size $c=n_{\mbox{\scriptsize RWE}}/n_{\mbox{\scriptsize RCT}}$ (replication to original), if we assume that the unit variance in both studies is the same. This implies that a replication study with a large sample size will show a smaller degree of replication success than a replication study with the same replication $p$-value, but a smaller sample size. The method hence requires replication studies with large sample sizes to be more convincing than those with small sample sizes to achieve the same degree of replication success.
@@ 3.1.2 Exact Type-I error control
The sceptical $p$-value has been recently re-calibrated to achieve exact overall Type-I error control (Micheloud et al.,, 2023) for every value of $c$. The new version has still improved project power compared to the two-trials rule but can now lead to replication success even if one of the two studies is not significant. In what follows we will report the re-calibrated "controlled" version of the sceptical $p$-value.
@@ 3.1.3 A comparison
The (controlled) sceptical $p$-value has exact overall Type-I error control (across both studies) at level $\alpha^{2}$ for all values of the variance ratio $c$. The two-trials rule also offers exact Type-I error control at level $\alpha^{2}$. However, the corresponding $p$-value $p_{\mbox{\scriptsize TTR}}$ cannot be smaller than $p_{\mbox{\scriptsize RCT}}$ nor $p_{\mbox{\scriptsize RWE}}$. Figure 2 compares the behaviour of the sceptical $p$-value $p_{S}$ and the corresponding $p$-value $p_{\mbox{\scriptsize TTR}}$ of the two-trials rule as a function of the relative sample size for fixed original $p$-values. The calculations assume that the relative effect size $\hat{\theta}_{\mbox{\scriptsize RWE}}/\hat{\theta}_{\mbox{\scriptsize RCT}}$ is fixed at 1 (top) and 0.75 (bottom), respectively. Note that for fixed relative effect size, the replication $p$-value depends on the assumed relative effect size and the relative sample size.
Figure 2 shows that the sceptical $p$-value behaves very similar to the two-trials rule if the relative sample size is small. However, it decreases smoothly towards zero, so can get smaller than $p_{\mbox{\scriptsize RCT}}$ if the relative sample size $c$ becomes large, whereas the two-trials rule $p$-value has an edge and then stays constant, as it cannot get smaller than $p_{\mbox{\scriptsize RCT}}$.
@@ 3.2 Power of the replication study
The power of the replication study, i.e. the probability that a replication study with a certain sample size leads to replication success, has been extensively studied (see for example Anderson and Maxwell,, 2017; Anderson and Kelley,, 2022; van Zwet and Goodman,, 2022; Micheloud and Held,, 2022; Micheloud et al.,, 2023). The power is usually calculated as the probability to reach a significant replication result conditional on the effect estimate from the original trial. This concept is illustrated in Figure 3 using the TRITON-TIMI trial from the RCT DUPLICATE initiative. The conditional power of the two-trials rule at the one-sided level $\alpha=0.025$ is plotted as a function of the effect estimate from the RCT (log hazard ratio) assuming that the variance is the same in the original trial and the emulation study ($c=1$). The original effect estimate in the TRITON-TIMI example is $\hat{\theta}_{\mbox{\scriptsize RCT}}=-0.21$, so with $c=1$, a conditional power of 98.1% is achieved. However, the original effect estimate comes with uncertainty, represented by the density in the top axis, and the 95% confidence interval $[-0.31,-0.11]$ around the original effect estimate. The conditional power for the values inside this confidence interval greatly vary, from 52.6% to 100%. One way to incorporate this uncertainty is to compute the predictive power (Spiegelhalter and Freedman,, 1986) by averaging the conditional power with respect to the distribution of the original effect estimate. The density of the conditional power is shown in the right axis of Figure 3 and the average is 92.6%, so considerably smaller than 98.1%.
@@
Figure 3: Conditional power as a function of the original effect estimate. The original effect estimate $\hat{\theta}_{\mbox{\scriptsize RCT}}=-0.21$ of the TRITON-TIMI trial and the corresponding conditional power of 98.1% are colored in red. The grey lines in the top and right axis are the distributions of the original effect estimate and the corresponding conditional power, respectively. The average conditional power of 92.6% (in blue) is the predictive power.
@@
If the power to detect the effect from the original study is not too large (< 95%), the conditional Type-I error rate (the probability of replication success if the replication effect estimate is zero) of the sceptical $p$-value is shown to be bounded and never larger than $2\alpha$ (Micheloud et al.,, 2023, Section 3.4). Such a two-fold increase in Type-I error rate is the price to be paid for increased power, but considered to be acceptable even in the regulatory setting (Rosenkranz,, 2002).
@@
Power calculations are also used to determine which sample size is required in the replication study to reach a certain power, usually between 80% and 90%. While replication sample size calculation is not relevant in the RCT DUPLICATE study, where all available patient information in the claims data was used, it is of interest to know the power of a RWE study with the available data. In Section 4, we calculate the conditional and predictive power of the two-trials rule and of the sceptical $p$-value for all the RWE emulations (Micheloud and Held,, 2022; Micheloud et al.,, 2023).
@@ 4.1 Replication of RCTs
Figure 4 shows $\hat{\theta}_{\mbox{\small RCT}}$ log(margin) versus $\hat{\theta}_{\mbox{\small RWE}}$ log(margin) for the 29 study pairs in the RCT DUPLICATE initiative. The margin is $1$ in the superiority studies. Points below the diagonal represent cases where the effect estimate from the RWE study is shrunken as compared to the effect estimate from the original trial. This shrinkage is towards 00 for superiority studies and towards the margin for non-inferiority studies. Replication effect estimates usually present a large amount of shrinkage as compared to their original counterpart in recent large-scale replication projects (e.g. in Open Science Collaboration,, 2015; Camerer et al.,, 2016, 2018). This is not the case here.
@@ 4.2 Replication power
Results are displayed in Table 2. The average conditional power (Amrhein et al.,, 2019; McShane et al.,, 2020) is 85.9% with the two-trials rule and 87.0% with the sceptical $p$-value. These values reduce to 83.5%, and 85.0%, respectively, for predictive power. If the true underlying effect is the same in the original trial and in the emulation study and both were conducted with high standards, we would expect the average predictive power to be equal to the proportion of success, which is smaller in the RCT DUPLICATE study (69% with both methods). The proportion of replication success as well as the average predictive power in the RCT DUPLICATE initiative for other values of $\alpha$ are shown in Figure A.1.
@@
For the 26 other studies, the power of the two-trials rule and the sceptical $p$-values is very similar, with slightly higher values with the sceptical $p$-values. This is because the original $p$-values are relatively small in every case, which is favored by the sceptical $p$-values (see Held et al., 2022b, ; Micheloud et al.,, 2023, Sections 3.1 and  3.4, respectively).
Cond. power Pred. power
Study $p_{RCT}$ two-trials rule sceptical $p$ two-trials rule sceptical $p$
@@
Table 2: Conditional and predictive power (in %) at the one-sided level $\alpha=0.025$ with both methods.
@@ 5 Discussion
A target trial emulation is a conceptual emulation of a "hypothetical" RCT, using a two-step process (Hernán and Robins,, 2016; Hernán et al.,, 2022). In the first step, the causal questions are defined in a protocol for the hypothetical RCT, which has to specify the key elements to define the causal estimands (i.e. eligibility criteria, treatment strategies, treatment assignment, the start and end of follow-up, outcomes, causal contrasts) and the data analysis plan (Hernán and Robins,, 2016). The RCT described in the protocol is defined as the target study for causal inference, which will be emulated using observational data in the second step (Hernán and Robins,, 2016; Hernán et al.,, 2022). Furthermore, several studies used an existing RCT instead of a hypothetical RCT, which will be emulated by the target trial emulation (Deng et al.,, 2022; Merola et al.,, 2022; Leening and Boersma,, 2024; Walker et al.,, 2024; Petito et al.,, 2020; McGrath et al.,, 2021; Khosrow-Khavar et al.,, 2022; Gallivan et al.,, 2023; Yoon et al.,, 2023). However, even with the existing RCT as a template for the target trial emulation, the effect estimates of the RWD emulation are not always compared with those of the original trial (Deng et al.,, 2022; Petito et al.,, 2020; McGrath et al.,, 2021; Khosrow-Khavar et al.,, 2022; Gallivan et al.,, 2023). A statistical comparison of the emulation with the original RCT further substantiates the results and increase the resulting evidence. As seen in the data presented, the sceptical $p$-value can be an excellent extension to evaluate the agreement and thus complements the evidence from target trial emulations of existing RCTs.
@@
The emulations of the 32 evaluated RCTs were previous implemented and analyzed by the RCT DUPLICATE Initiative (Wang et al.,, 2023). In this project, Wang et al., (2023) used different binary agreement metrics to check the replication success of the 32 emulations using claims data of three different US insurance companies. First, named full statistical significance agreement, is fulfilled, if the estimates (log($HR$)) and the related 95% CIs are on the same side of the null (without consideration of the predefined non-inferiority margins in the case of non-inferiority trials). 23 (72%) of the 32 emulations met statistical significance agreement (Wang et al.,, 2023). This binary agreement metric should be equivalent to the two-trials rule. However, for two studies, i.e. D5896 and PRONOUNCE, the $p$-value of the two-trials rule delivered a different conclusion of the replication success. Wang et al., (2023) reported that statistical agreement was fulfilled for all of the two emulations, whereby in our analysis, $p_{\mbox{\scriptsize TTR}}>0.025$ holds. In the case of PRONOUNCE, the original RCT had a non-significant treatment effect, resulting in an non-significant $p_{\mbox{\scriptsize TTR}}$ due to the definition of the $p$-value. Moreover, the formalism of $p_{\mbox{\scriptsize TTR}}$ is not meaningful in the case of non-significant effects of the original trial. In the case of the D5896 trial, Wang et al., (2023) did not use the non-inferiority margins for the definition of full statistical agreement, as it was simply defined as having point estimates and confidence intervals on the same side of null. Using sceptical p-value, it is possible to account for the non-inferiority margins, which is one major advantage of this formalism compared to the binary agreement metrics.
@@
As the second binary agreement metric, Wang et al., (2023) analysed estimated agreement, whereby the estimate of the emulation has to be an element of the 95% CI of the original effect, which was met for 66% of the emulations. As the last one, Wang et al., (2023) evaluated, if the absolute standardized differences of the estimates are smaller then 1.96, which was fulfilled in 72% of the emulations. Using the sceptical $p$-value, 69% of the emulations were evaluated as successful replications. Beside the emulations with non-significant original effects, the findings of the study at hand are in accordance to findings of Wang et al., (2023).
@@
However, in general, it is not always possible to replicate the results of an RCT, especially if there are fundamental design differences between the RCT and RWE (Wang et al.,, 2023; Heyard et al.,, 2024). Wang et al., (2023) stated that the conceptual emulation of the original RCT leads to an independent study that emulate the design of the original trial instead of replicating the population of the RCT. Potential differences of the two population (e.g. different distribution of comorbidities) may result in differences in the effect size estimates (Wang et al.,, 2023). Moreover, Heyard et al., (2024) analyzed differences between the original RCTs and the related RWE emulations using meta-regression models. They concluded that most of the heterogeneity between RCT and RWD emulation could be explained by delayed effect of treatment, discontinuation of treatment during run-in period, and treatment started in hospital (Heyard et al.,, 2024).
@@
There are other studies available that emulate RCTs using RWD, where the binary metrics regulatory agreement, estimated agreement and standardized difference agreement were often used to evaluate the agreement of the effect estimates from the original trials and the related emulation study (Jang et al.,, 2022; Jin et al.,, 2022; Merola et al.,, 2023; Antoine et al.,, 2024; Signori et al.,, 2024; D’Andrea et al.,, 2024). However, the two-trials rule and the sceptical $p$-values are good measures to evaluate the agreement between an original RCT and the corresponding emulation study and have benefits compared to binary metrics. Besides the advantages disscussed above, both methods allow to calculate the power of the emulation study based on the effect estimate found in the RCT, with and without uncertainty incorporation. For all 29 studies considered here, the conditional and predictive power is larger with the sceptical $p$-values than with the two-trials rule. Furthermore, the replication rate is in general slightly higher with the sceptical $p$-value.
@@
In this paper we focused on the controlled version of the sceptical $p$-value, to ensure a fair comparison with the two-trials rule in terms of overall Type-I error control. An alternative to the controlled re-calibration is the golden re-calibration of the sceptical $p$-value (Held et al., 2022b, ). The golden version is less stringent than the original formulation, because, to establish replication success, original trial and replication study do not both necessarily need to be significant at level $\alpha$. However, a borderline significant original trial ($p_{\mbox{\scriptsize RCT}}=\alpha$) cannot lead to replication success, if there is shrinkage of the replication effect estimate. As shrinkage of the effect estimate from the emulation study as compared to effect estimate from the original trial seems to not be an issue in this context, one could favor the controlled sceptical $p$-value over the golden sceptical $p$-value. Moreover, a comparison of the golden version with the two-trials rule is more difficult, as the two methods have different Type-I error rates.
@@
The presented work had several limitations. No formalism to assess the replication success of an original null effect using the sceptical $p$-value was available (Pawel et al.,, 2024). Recently Micheloud and Held, (2024) provided a formalism of the sceptical $p$-value for equivalence studies. However, a trial that was initially planned for superiority, but fails to show a significant treatment effect, should not be interpreted as an equivalence study and, thus, the formalism for equivalence studies should not be assigned in this case. Moreover, as stated above, fundamental design differences between RCT and RWD emulation may result in bias, which will also affect the effect estimate of the RWD study.

### v2

@@ 3.1.1 The original formulation
The sceptical $p$-value can be interpreted quantitatively: Smaller $p_{S}$ values indicate a higher degree of replication success. The sceptical $p$-value is one-sided to ensure that replication can only occur if the replication effect estimate has the same sign as the original effect estimate. Further properties of the sceptical $p$-value are studied in Held (2020). First, the approach can be shown to be more stringent than the two-trials rule: if replication success is achieved ($p_{S}\leq\alpha$), then two-trials rule is also fulfilled ($p_{\mbox{\scriptsize TTR}}=\max\{p_{\mbox{\scriptsize RCT}},p_{\mbox{\scriptsize RWE}}\}\leq\alpha$) but not necessarily the other way around. As a consequence, the overall Type-I error rate of the sceptical $p$-value (assuming that both the original and the replication effect estimate is zero) is smaller than the Type-I error rate $\alpha^{2}$ of the two-trials rule. Second, for fixed original and replication $p$-value, the sceptical $p$-value increases with increasing variance ratio $c$. The variance ratio can be rewritten as the relative sample size $c=n_{\mbox{\scriptsize RWE}}/n_{\mbox{\scriptsize RCT}}$ (replication to original), if we assume that the unit variance in both studies is the same. This implies that a replication study with a large sample size will show a smaller degree of replication success than a replication study with the same replication $p$-value, but a smaller sample size. The method hence requires replication studies with large sample sizes to be more convincing than those with small sample sizes to achieve the same degree of replication success.
@@ 3.1.2 Exact Type-I error control
The sceptical $p$-value has been recently re-calibrated to achieve exact overall Type-I error control across both studies (Micheloud et al., 2023) for every value of $c$. The new version has still improved project power compared to the two-trials rule but can now lead to replication success even if one of the two studies is not significant. In what follows we will report the re-calibrated "controlled" version of the sceptical $p$-value.
The exact overall calibration at level $\alpha^{2}$ implies that we can invert the sceptical $p$-value to obtain one-sided confidence intervals for the combined effect estimate. The upper limit $\theta_{u}$ of a 97.5% CI for the log hazard ratio is therefore obtained as the value of $\delta=\theta_{u}$, where the modified test statistic (1) gives a squared sceptical $p$-value of 0.025. Exponentiation gives an upper limit on the hazard ratio scale, which can directly be compared to the non-inferiority (or superiority) margin. This approach provides an alternative to standard meta-analytic pooling, which may not be compatible to the sceptical $p$-value nor the two-trials rule.
@@ 3.1.3 A comparison
The (controlled) sceptical $p$-value has exact overall Type-I error control (across both studies) at level $\alpha^{2}$ for all values of the variance ratio $c$. The two-trials rule also offers exact Type-I error control at level $\alpha^{2}$. However, the corresponding $p$-value $p_{\mbox{\scriptsize TTR}}=\max\{p_{\mbox{\scriptsize RCT}},p_{\mbox{\scriptsize RWE}}\}$ cannot be smaller than $p_{\mbox{\scriptsize RCT}}$ nor $p_{\mbox{\scriptsize RWE}}$. Figure 2 compares the behaviour of the sceptical $p$-value $p_{S}$ and the corresponding $p$-value $p_{\mbox{\scriptsize TTR}}$ of the two-trials rule as a function of the relative sample size for fixed original $p$-values. The calculations assume that the relative effect size $\hat{\theta}_{\mbox{\scriptsize RWE}}/\hat{\theta}_{\mbox{\scriptsize RCT}}$ is fixed at 1 (top) and 0.75 (bottom), respectively. Note that for fixed relative effect size, the replication $p$-value depends on the assumed relative effect size and the relative sample size.
Figure 2 shows that the sceptical $p$-value behaves very similar to the two-trials rule if the relative sample size is small. However, it decreases smoothly towards zero and can get smaller than $p_{\mbox{\scriptsize RCT}}$ if the relative sample size $c$ becomes large, whereas the two-trials rule $p$-value has an edge and then stays constant, as it cannot get smaller than $p_{\mbox{\scriptsize RCT}}$.
@@ 3.2 Power of the replication study
The power of the replication study, i.e. the probability that a replication study with a certain sample size leads to replication success, has been extensively studied (see for example Anderson and Maxwell, 2017; Anderson and Kelley, 2022; van Zwet and Goodman, 2022; Micheloud and Held, 2022; Micheloud et al., 2023). The power is usually calculated as the probability to reach a significant replication result conditional on the effect estimate from the original trial. This concept is illustrated in Figure 3 using the TRITON-TIMI trial from the RCT DUPLICATE initiative. The conditional power of the two-trials rule at the one-sided level $\alpha=0.025$ is plotted as a function of the effect estimate from the RCT (log hazard ratio) assuming that the variance is the same in the original trial and the emulation study ($c=1$). The original effect estimate in the TRITON-TIMI example is $\hat{\theta}_{\mbox{\scriptsize RCT}}=-0.21$, so with $c=1$, a conditional power of 98.1% is achieved. However, the original effect estimate comes with uncertainty, represented by the density in the top axis, and the 95% confidence interval $[-0.31,-0.11]$ around the original effect estimate. The conditional power for the values inside this confidence interval greatly vary, from 52.6% to 100%. One way to incorporate this uncertainty is to compute the predictive power (Spiegelhalter and Freedman, 1986) by averaging the conditional power with respect to the distribution of the original effect estimate. The density of the conditional power is shown in the right axis of Figure 3 and the average is 92.6%, so considerably smaller than 98.1%.
@@
Figure 3: Conditional power as a function of the original effect estimate. The original effect estimate $\hat{\theta}_{\mbox{\scriptsize RCT}}=-0.21$ of the TRITON-TIMI trial and the corresponding conditional power of 98.1% are colored in red. The grey lines in the top and right axis are the distributions of the original effect estimate and the corresponding conditional power, respectively. The average conditional power of 92.6% (in blue) is the predictive power. The green intervals represent the range given by the 95% confidence interval of the original effect estimate.
@@
If the power to detect the effect from the original study is not too large (< 95%), the conditional Type-I error rate (the probability of replication success if the replication effect estimate is zero) of the sceptical $p$-value is shown to be bounded and never larger than $2\alpha$ (Micheloud et al., 2023, Section 3.4). Such a two-fold increase in Type-I error rate is the price to be paid for increased power, but considered to be acceptable even in the regulatory setting (Rosenkranz, 2002).
@@
Power calculations are also used to determine which sample size is required in the replication study to reach a certain power, usually between 80% and 90%. While replication sample size calculation is not relevant in the RCT DUPLICATE study, where all available patient information in the claims data was used, it is of interest to know the power of a RWE study with the available data. It is therefore advisable to determine the power for an existing dataset before calculating the effect estimates for the emulation, to ensure that the data are indeed sufficient. Conversely, calculating the minimum necessary sample size can also be beneficial in the context of RWE, for example, to understand the cohort size required after propensity score matching. If the power is insufficient or if the sample size is inadequate after matching, the chosen dataset is likely not suitable for the planned emulation. In Section 4, we calculate the conditional and predictive power of the two-trials rule and of the sceptical $p$-value for all the RWE emulations (Micheloud and Held, 2022; Micheloud et al., 2023).
@@ 4.1 Replication of RCTs
Figure 4 shows $\hat{\theta}_{\mbox{\small RCT}}$ log(margin) versus $\hat{\theta}_{\mbox{\small RWE}}$ log(margin) for the 29 study pairs in the RCT DUPLICATE initiative. The margin is $1$ in the superiority studies. Points below the diagonal represent cases where the effect estimate from the RWE study is shrunken as compared to the effect estimate from the original trial. This shrinkage is towards 00 for superiority studies and towards the margin for non-inferiority studies. Replication effect estimates usually present a large amount of shrinkage as compared to their original counterpart in recent large-scale replication projects (e.g. in Open Science Collaboration, 2015; Camerer et al., 2016, 2018). This is not the case here.
@@ 4.2 Replication power
Results are displayed in Table 2. The average conditional power (Amrhein et al., 2019; McShane et al., 2020) is 85.9% with the two-trials rule and 87.0% with the sceptical $p$-value. These values reduce to 83.5%, and 85.0%, respectively, for predictive power. If the true underlying effect is the same in the original trial and in the emulation study and both were conducted with high standards, we would expect the average predictive power to be equal to the proportion of success, which is smaller in the RCT DUPLICATE study (69% with both methods). The proportion of replication success as well as the average predictive power in the RCT DUPLICATE initiative for other values of $\alpha$ are shown in Figure A.1.
@@
For the 26 other studies, the power of the two-trials rule and the sceptical $p$-values is very similar, with slightly higher values with the sceptical $p$-values. This is because the original $p$-values are relatively small in every case, which is favored by the sceptical $p$-values (see Held et al., 2022; Micheloud et al., 2023, Sections 3.1 and  3.4, respectively).
4.3 Confidence intervals
We have also calculated confidence intervals for the hazard ratio based on inversion of the sceptical $p$-value and compared those to the ones obtained from applying a fixed effect meta-analysis to the results from the RCT and RWD. The results are given in Table B.1 in Appendix B. The upper limit of both confidence intervals are similar, the upper limit of the sceptical CI being often slightly larger than the meta-analytic one. The differences get larger when the uncertainty of the effect estimates is substantially different, for example for DAPA-CKD or INSPIRE, where the meta-analytic pooled effect estimate is dominated by the larger study. This is also the case for the three superiority studies PARADIGM-HF, IMPACT and POET-COPD, where the meta-analytic confidence interval indicates significance (at the standard two-sided 95% level) with an upper CI limit below 1, although the RWE effect estimate is above 1 in all three cases. In contrast, the upper limit of the sceptical CI is always above one, in line with the large sceptical $p$-values for those studies reported Table 1. This illustrates that meta-analytic pooling can still flag a positive (significant) finding, even if one of the two studies has an effect in the wrong direction (Held, 2024).
Conditional power Predictive power
Study $p_{\mbox{\scriptsize RCT}}$ two-trials rule sceptical $p$ two-trials rule sceptical $p$
@@ 4.3 Confidence intervals
Table 2: Conditional and predictive power (in %) at the one-sided level $\alpha=0.025$ with the two-trials rule and the sceptical $p$-value.
@@ 5 Discussion
A target trial emulation is a conceptual emulation of a "hypothetical" RCT, using a two-step process (Hernán and Robins, 2016; Hernán et al., 2022). In the first step, the causal questions are defined in a protocol for the hypothetical RCT, which has to specify the key elements to define the causal estimands (i.e. eligibility criteria, treatment strategies, treatment assignment, the start and end of follow-up, outcomes, causal contrasts) and the data analysis plan (Hernán and Robins, 2016). The RCT described in the protocol is defined as the target study for causal inference, which will be emulated using observational data in the second step (Hernán and Robins, 2016; Hernán et al., 2022). Furthermore, several studies used an existing RCT instead of a hypothetical RCT, which will be emulated by the target trial emulation (Deng et al., 2022; Merola et al., 2022; Leening and Boersma, 2024; Walker et al., 2024; Petito et al., 2020; McGrath et al., 2021; Khosrow-Khavar et al., 2022; Gallivan et al., 2023; Yoon et al., 2023). However, even with the existing RCT as a template for the target trial emulation, the effect estimates of the RWD emulation are not always compared with those of the original trial (Deng et al., 2022; Petito et al., 2020; McGrath et al., 2021; Khosrow-Khavar et al., 2022; Gallivan et al., 2023). A statistical comparison of the emulation with the original RCT further substantiates the results and increase the resulting evidence. As seen in the data presented, the sceptical $p$-value can be an excellent extension to evaluate the agreement and thus complements the evidence from target trial emulations of existing RCTs.
@@
The emulations of the 32 evaluated RCTs were previous implemented and analyzed by the RCT DUPLICATE Initiative (Wang et al., 2023). In this project, Wang et al. (2023) used different binary agreement metrics to check the replication success of the 32 emulations using claims data of three different US insurance companies. First, named full statistical significance agreement, is fulfilled, if the estimates (log(HR)) and the related 95% CIs are on the same side of the null (without consideration of the predefined non-inferiority margins in the case of non-inferiority trials). 23 (72%) of the 32 emulations met statistical significance agreement (Wang et al., 2023). This binary agreement metric should be equivalent to the two-trials rule. However, for two studies, i.e. D5896 and PRONOUNCE, the $p$-value of the two-trials rule delivered a different conclusion of the replication success. Wang et al. (2023) reported that statistical agreement was fulfilled for all of the two emulations, whereby in our analysis, $p_{\mbox{\scriptsize TTR}}>0.025$ holds. In the case of PRONOUNCE, the original RCT had a non-significant treatment effect, resulting in an non-significant $p_{\mbox{\scriptsize TTR}}$ due to the definition of the $p$-value. Moreover, the formalism of $p_{\mbox{\scriptsize TTR}}$ is not meaningful in the case of non-significant effects of the original trial. In the case of the D5896 trial, Wang et al. (2023) did not use the non-inferiority margins for the definition of full statistical agreement, as it was simply defined as having point estimates and confidence intervals on the same side of null. Using sceptical $p$-value, it is possible to account for the non-inferiority margins, which is one major advantage of this formalism compared to the binary agreement metrics.
@@
As the second binary agreement metric, Wang et al. (2023) analysed estimated agreement, whereby the estimate of the emulation has to be an element of the 95% CI of the original effect, which was met for 66% of the emulations. As the last one, Wang et al. (2023) evaluated, if the absolute standardized differences of the estimates are smaller then 1.96, which was fulfilled in 72% of the emulations. Using the sceptical $p$-value, 69% of the emulations were evaluated as successful replications. Beside the emulations with non-significant original effects, the findings of the study at hand are in accordance to findings of Wang et al. (2023).
@@
However, in general, it is not always possible to replicate the results of an RCT, especially if there are fundamental design differences between the RCT and RWE (Wang et al., 2023; Heyard et al., 2024). Wang et al. (2023) stated that the conceptual emulation of the original RCT leads to an independent study that emulate the design of the original trial instead of replicating the population of the RCT. Potential differences of the two population (e.g. different distribution of comorbidities) may result in differences in the effect size estimates (Wang et al., 2023). Moreover, Heyard et al. (2024) analyzed differences between the original RCTs and the related RWE emulations using meta-regression models. They concluded that most of the heterogeneity between RCT and RWD emulation could be explained by delayed effect of treatment, discontinuation of treatment during run-in period, and treatment started in hospital (Heyard et al., 2024). The sceptical $p$-value itself cannot distinguish between replication failure due to design differences, incorrect model specification or differences due to population differences. Instead, the sceptical $p$-value evaluates and quantifies the agreement between two effect estimators. A successful replication of an RCT using RWD, indicated by a significant sceptical $p$-value, suggests the external validity of the RCT results. Conversely, if the replication is unsuccessful, the p-value does not provide insight into the reasons. These issues must be addressed using other methods. In studies based on RWD, it is essential to adequately address data heterogeneity methodologically, for example, using propensity score methods. An important aspect is achieving sufficient balance between the treatment groups. However, it is possible that the treatment effect observed in an RCT cannot be replicated in RWD due to fundamentally different cohorts. This is less a failure of RWE and may indicate that external validity is limited.
@@
There are other studies available that emulate RCTs using RWD, where the binary metrics regulatory agreement, estimated agreement and standardized difference agreement were often used to evaluate the agreement of the effect estimates from the original trials and the related emulation study (Jang et al., 2022; Jin et al., 2022; Merola et al., 2023; Antoine et al., 2024; Signori et al., 2024; D’Andrea et al., 2024). However, the two-trials rule and the sceptical $p$-values are good measures to evaluate the agreement between an original RCT and the corresponding emulation study and have benefits compared to binary metrics. Besides the advantages discussed above, both methods allow to calculate the power of the emulation study based on the effect estimate found in the RCT, with and without uncertainty incorporation. For all 29 studies considered here, the conditional and predictive power is larger with the sceptical $p$-values than with the two-trials rule. Furthermore, the replication rate is in general slightly higher with the sceptical $p$-value.
@@
In this paper, we focused on the controlled version of the sceptical $p$-value, to ensure a fair comparison with the two-trials rule in terms of overall Type-I error control. An alternative to the controlled re-calibration is the golden re-calibration of the sceptical $p$-value (Held et al., 2022). The golden version is less stringent than the original formulation, because, to establish replication success, original trial and replication study do not both necessarily need to be significant at level $\alpha$. However, a borderline significant original trial ($p_{\mbox{\scriptsize RCT}}=\alpha$) cannot lead to replication success, if there is shrinkage of the replication effect estimate. As shrinkage of the effect estimate from the emulation study as compared to effect estimate from the original trial seems to not be an issue in this context, one could favor the controlled sceptical $p$-value over the golden sceptical $p$-value. Moreover, a comparison of the golden version with the two-trials rule is more difficult, as the two methods have different Type-I error rates.
@@
The presented work had several limitations. No formalism to assess the replication success of an original null effect using the sceptical $p$-value was available (Pawel et al., 2024). Recently Micheloud and Held (2024) provided an extension of the sceptical $p$-value to equivalence studies. However, a trial that was initially planned for superiority, but fails to show a significant treatment effect, should not be interpreted as an equivalence study and, thus, the formalism for equivalence studies should not be assigned in this case. Instead of performing a second study with the wrong design, a new study with the aim of proofing the equivalence should be initialized (and then replicated). Moreover, as stated above, fundamental design differences between RCT and RWD emulation may result in bias, which will also affect the effect estimate of the RWD study.